OpenStudy (albert0898):

If the average of q, r, s, and t is 10, what is the average of s - r, 3r + t, 10 - r, and 6 + q?

myininaya (myininaya):

\[\frac{q+r+s+t}{4}=10\] We want to know the average of \[\frac{(s-r)+(3r+t)+(10-r)+(6+q)}{4}=?\] did you try combining like terms :)

OpenStudy (albert0898):

OHHHHHHHHH. This is the moment where I realize how easy this is, and call myself an idiot.

myininaya (myininaya):

omg don't call yourself an idiot lol

OpenStudy (albert0898):

:3